OpenStudy (anonymous):

30 mi / h = ____ ft / min

OpenStudy (anonymous):

2640

OpenStudy (anonymous):

\[\frac{30\text{ miles}}{1\text{ hour}}\times\frac{1\text{ hour}}{60\text{ min}}\times\frac{5280\text{ ft}}{1\text{ mile}}=\underline{?}\frac{\text{ft}}{\text{min}}\]